OpenStudy (anonymous):

Which of the following is an irrational number? (10square root 2) 4__+__ square root 8 0.___ 63 square root of 196 ???

jimthompson5910 (jim_thompson5910):

not sure what the underscores are for

OpenStudy (anonymous):

Oh, sorry they it's underlined so whatever that means.

jimthompson5910 (jim_thompson5910):

but \(\large \sqrt{2}\) is an irrational number so is \(\large 10\sqrt{2}\)

jimthompson5910 (jim_thompson5910):

\(\large \sqrt{8}\) is an irrational number, and so is \(\large 4+\sqrt{8}\)

jimthompson5910 (jim_thompson5910):

0.63 is rational since 0.63 = 63/100

jimthompson5910 (jim_thompson5910):

\(\large \sqrt{196} = 14\), so that's a rational number

OpenStudy (anonymous):

So...my awnser would be 4+ square root of 8?

jimthompson5910 (jim_thompson5910):

that's irrational, but so is \(\large 10\sqrt{2}\)

jimthompson5910 (jim_thompson5910):

are there any exponents over the \(\large 10\sqrt{2}\)

OpenStudy (anonymous):

Okay, that makes sense. Thanks.

jimthompson5910 (jim_thompson5910):

if not, then both A and B are irrational